OpenStudy (anonymous):

The molar mass of a certain gas is 58.1 g/mol. What is the density of the gas in g/L at STP?

OpenStudy (aaronq):

use the ideal gas law, PV=nRT -> P=nRT/V sub in: n=m/M P=mRT/MV sub in: d=m/V P=dRT/M rearrange.. d=PM/RT

OpenStudy (anonymous):

where little m is mass in grams and big M is molar mass.